Problem Four

Luke Rucker is married with two withholdings and an hourly employee at ABC Widgets. The company pays its employees on a biweekly basis. Luke earns $20 per hour. For the July 31, 20XX pay, Luke worked 92 hours.

Required:

Assuming a 40-hour week, complete Luke’s payroll register below. Round all answers to two decimal places. Complete the payroll register with the period’s information assuming 40-hour week.

Employee Name                                                                                                                             

                                                                                                                               

Company                                                                                                                           

                                                                                                                               

Marital Status   

# of Withholdings            

Hourly Pay Rate Overtime Pay Rate          

# of Reg Hours  

# of OT Hrs         

Reg Earn             

OT Earn Gross Earnings
Answer #3

Employee Name: Luke Rucker

Company :- ABC Widgets

Marital status: Married

# of withholdings: 2

Hourly Wage Rate - 20$

Overtime Pay rate = 20*150% as per US Federal Law.

= 30$

Regular Earnings = 20 * 80 = 1600$

Overtime Earnings =( 92-80)* 30 = 360$

Gross Earnings b= 1960$


answered by: Eyesst
Answer #1

Weekly earning=40hrs*20=800$

Total hrs worked =90

Extra hrs worked =90-40 =50

=50*20

=1000$

Total earnings =800$+ 1000$

=1800$
Answer #2

Married 

 2

20 

 30

 92

52 

800 

1560 

2360 


source: ME
answered by: ANNA